The data below represents the number of complaints recorded the customer service desk at a local call center. The data shows the number of complaints logged each day over a seven day period.

Day

Complaints Received

1

12

2

2

3

7

4

1

5

3

6

0

7

1

  1. What is the upper control limit of the c-chart?
    1. 9.50
    2. 3.45
    3. 12.00
    4. Doesn’t exist

  1. Which of the following are true about the process?
    1. Special causes of variation are indicated
    2. There are only common causes of variation
    3. The process is in control
    4. Process is within specification limits.
    5. B, C and D are correct
Answer #1
total defects d = 26.00
total samples m= 7.00
average number of defects per sample c=d/m= 3.71
control line (CL) = c= 3.71
upper control limit(UCL) =c+3*(√c)= 9.50
lower control limit(LCL) =max(0,c-3*(√c))= 0.00

1)

upper control limit of the c-chart =9.50 option A

2) Special causes of variation are indicated   : option A: as 12 is above upper control limit